OpenStudy (anonymous):

Im having so much trouble with this question, someone help pls :( For each function shown: Name the starting point. Explain how y changes as x increases. Function #1: y = 0.5 + 3x Function #2: y = 3(0.5)^x

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Let's begin with function #1 y = 0.5 + 3x The question asks "what is the 'starting' point"? Think about that. What does it mean to start at a point? You need to "assume" that x starts at some number, say the numerical 0 is a good choice. So, now replace x with 0 in the defining equation of the function to give you the "starting value" of y. y = 0.5 + 3x y = 0.5 + 3•(0) y = 0.5 + 0 y = 0.5 So, this gives you the "starting point" (x, y) = (0, 0.5).

OpenStudy (skullpatrol):

Any questions? :-)

OpenStudy (skullpatrol):

Use the exact same procedure to find the starting point for function #2

OpenStudy (skullpatrol):

The second part of the question asks for an explanation of how y changes as x increases. To explain this it's best to draw a graph. But where do we start? Let's use the starting point calculated from the first part of the question for y = 0.5 + 3x we calculated that (0, 0.5) was the starting point, let's replace x with the number 1 and see what happens to y. y = 0.5 + 3x y = 0.5 + 3•(1) y = 0.5 + 3 y = 3.5 So, now we have another point on the graph of y = 0.5 + 3x, namely (1, 3.5). Let's compare it to the starting point: (0, 0.5) ----> (1, 3.5) In other words, for every 1 unit increase in x, there must be a 3 unit increase in y because of the 3x term in the given linear function.

OpenStudy (skullpatrol):

Note: the same procedure can be used for function #2, but the given function is not linear ie of the form y = mx +b.